You are here

ENR CTIO Blanco 4-m 2014-11-30

Primary tabs

Date: 
Sunday, November 30, 2014
Telescope: 
CTIO Blanco 4-m
Submitter: 
Chris Smith
Submitter Email: 
Observer Support: 
Rodrigo Hernandez
Night Assistant: 
Claudio Aguilera
Also Present: 
Lucinda Clerkin, Nikolay Kuropatkin, Javier Rojas
Problems: 

None.

Comments: 
*Observing Team*
 
_DES Observers:_ Nikolay Kuropatkin (Obs1), Chris Smith (Obs2), Lucy Clerkin (RM) 
_Telops:_ Rodrigo Hernandez (Obs Support), Claudio Aguilera (Night Support)
_Remote Support:_ ---
 
*Observing Plan*
 
# Standard bias, standard flats
# evening standard stars 
# run ObsTac queue
# end of the night standard star
# end of night calibrations 
 
*Conditions*
 
# weather in afternoon 16:30 (local): clear sky, 13.0C , wind 15 mph, humidity 42%
# at sunset 20:30 local: 
# 22:00, after starting OCS: 9.7C, wind 11mph, humidity 32% 
# NOTE: mirror temp significantly lower than outside temp (about -2C)
# 11:15 local: oddly enough, temp drifted downward to 10.1, then jumped back up at around 11:15 to 10.8!  wind steady at around 10mph
# 12:15 local: temp drifting down again, 10.6C, wind 14mph, humidty 31%
# 02:30 local: temp stabilizing at 10.0C, wind 7mph, humidity 25%
# 04:10 local: temp 9.5C, wind 6 mph, humidity 30%
# 05:25 local: temp 10.1, wind 10 mph, humidity 35%
 
# weather summary: mostly clear, 9-13C, humidity 20-35%
# Seeing conditions: ~0.9 to 1.5, jumpy early, steady in middle, rising at end
# Wind conditions: wind 5-15 mph
 
*Accomplishments*
 
# Afternoon: SISPI restarted, zeros done, (YziVRrgu) dome flats done, flats checked OK
# Afternoon: SISPI restarted again (to pick up changes OCS put in by Eric)
# Dome opened around 19:30 local
# At 21:00 local (after sunset) 9.7C, wind 11.5mph, humidity 39% 
# NOTE: Windscreen up from BEGINNING of night
# Zero point Correction: RA Offset -8.2, Dec Offset -0.1
# Begin of night: Standards 
## SDSS J0320-0000
## SDSS J2140-0000
## SDSS J0100-0000
# 21:43 local: start ObsTac queue; SUPERNOVAE (last night's data not good)
# SN field E1 (griz), seeing g=1.12", r=1.00"
# NOTE: primary mirror still about 1deg COLDER than outside air (good!)
# SN field E2 (griz), seeing g=1.10", r=1.05"
# SN field X2 (griz), seeing g=1.23", r=1.09", z=1.03"
# SN field C3 (z), seeing z=1.00"
# SN field C2 (griz), seeing g=1.32", z=1.14"
# SN field C1 (griz), seeing g=1.39", z=0.91"
# NOTE: primary mirror cooling turned off at 12:15 local (image 383146)
# 00:35 local: CHANGE TO WIDE-FIELD SURVEY observations
# start in Y band (due to bright moon), mostly Y with some z mixed in
# seeing around 0.9" to 1.0" for most of Y and z images
# as Moon sets, sky still completely clear!
# survey exposures: i/r/g bands starting at 03:00 local
# moon set at 03:08 local
# End of night: Standards (05:25 local)
## SDSS J1048-0000
## SDSS J0843-0000  
 
 
*Problems tagged as Synopsis_Problem*
 
# Rasicam: off
# DIMM: on
 
 
*Minor Problems and Relevant Information*
 
# primary mirror cooling time extended into night;
# cooling turned off at 12:15 local, 
# before turning off mirror cooling (at 12:15), 
## mirror (pmas_temp) was 1.5 degrees COLDER than outside_temp, and 2.5 degrees COLDER than lowdome_temp
# one hour after turning off mirror cooling (1:15)
## mirror (pmas_temp) rose to EQUAL outside_temp, still about 1 degree colder than lowdome
# two hours after turnning off mirror cooling (2:15)
## mirror (pmas_temp) still rising, now about 0.7 degree warmer than outside_temp, but still about 0.5 degrees colder than lowdome_temp.
# see plot in E-log!
 
*Notes for Data Management*
 
# clear, variable seeing from 0.9 to 1.4"
# data should be good! 
 
Conditions: 
General Conditions
Comment: 
None.
1st Quarter
Clouds: 
Clear
Seeing Value: 
1.25
Seeing Variability: 
0.2
Seeing Source: 
The Telescope I used
Sky : 
0
2nd Quarter
Clouds: 
Clear
Seeing Value: 
1.20
Seeing Variability: 
0.2
Seeing Source: 
The Telescope I used
Sky: 
0
3rd Quarter
Clouds: 
Clear
Seeing Value: 
0.90
Seeing Variability: 
0
Seeing Source: 
The Telescope I used
Sky: 
0
4th Quarter
Clouds: 
Clear
Seeing Value: 
1.10
Seeing Variability: 
0.2
Seeing Source: 
The Telescope I used
Sky: 
0
Program: 
Proposal: 
2012B-0001
Target of Oportunity: 
No
Time Observed: 
8.0
Time Engineering: 
0.0
Time lost technical: 
0.0
Time lost weather: 
0.0
Time lost other: 
0.0
Description: 
Dark Energy Survey
Primary Investigator: 
Frieman
Institution: 
DES
Instrument: 
CTIO 4m DECam
Organization: 
NOAO
Total Program Time: 
8
Total Times Time Observed Time Engineering Time lost technical Time lost weather Time lost other Total Program Time
2014-11-30 8.0 0.0 0.0 0.0 0.0 8